A large number of applicants for graduate studies in business administration are administered an aptitude test. The scores are normally distributed with an average of 460 and a standard deviation of 80. What fraction of the applicants would expect to have scores of 600 or more?

a. 0.4599

b. 0.5401

c. 0.0852

d. 0.0401
Answer #1

Answer:

Given,

To determine that which fraction of the applicants would expect to have scores of 600 or more

Standard deviation = 80

mean = 460

Now consider,

P(X > 600) = P(Z > (xbar - mean)/standard deviation)

substitute values

P(X > 600) = P(Z > (600 - 460) / 80)

= P(Z > 140/80)

= P(Z > 1.75)

= 0.0400592

= 0.0401

So option D is correct answer.